A fast cleaner clean 10 rooms in an hour and a slower cleaner 4 rooms in an hour if both work together how long will take to clean 26 rooms
Found 2 solutions by ptaylor, Alan3354:
Answer by ptaylor(2198)

Let x=time it takes if both working together to clean 26 rooms
Both working together cleans at the rate of 26/x rooms/hour
Fast cleaner cleans at the rate of 10 rooms/hour
Slower cleaner cleans at the rate of 4 rooms/hour
Our equation to solve then is:
10+4=(26/x);14=26/x multiply each term by x
14x=26
x=26/14=13/7 =1.9 hours
CHECK
(I like to deal in fractions rather than decimals; less messy and eliminates round off errors, sooo:)
In (13/7)) hours, fast cleaner cleans (13/7)*10=130/7 rooms
In (13/7) hours, slower cleaner cleans (13/7)*4=52/7 rooms
(130/7)+(52/7) should equal 26
182/7=26
26=26
Hope this helps----ptaylor

Answer by Alan3354(69443)

A fast cleaner clean 10 rooms in an hour and a slower cleaner 4 rooms in an hour if both work together how long will take to clean 26 rooms
-----------
Together, they do 14 rooms per hour.
Together it takes 26/14 hours.